OpenStudy (anonymous):

what is the value of m: 4+m/8=3/4

OpenStudy (anonymous):

subtract 4 from both sides 4 = 16/4

OpenStudy (anonymous):

so m equals 19/4 ?

OpenStudy (anonymous):

I don't know m/8=3/4 -16/4 m/8 = -13/4

OpenStudy (anonymous):

o ok thanks ill sthill give you a medal though

OpenStudy (anonymous):

m = -80/4 - 24/4 m = -104/4 ???

OpenStudy (anonymous):

thanks